A closed coil of 40 turns and of area 200 cm2, is rotated in a magnetic field of flux density 2 Wb m-2. it rotates from a position where its plane makes an angle of 30° with the field to a position perpendicular to the field in atime 0.2 sec. Find the magnitude of the emf induced in the coil due to its rotation.


1
Expert's answer
2022-03-10T10:17:35-0500

Given:

"\Delta \theta=\pi/3"

"\Delta t=0.2\:\rm s"

"N=40"

"A=200*10^{-4}\:\rm m^2"

"B=2\:\rm T"


The magnitude of the EMF induced in the coil

"{\cal E}=NBA\omega=NBA\frac{\Delta \theta}{\Delta t}"

"{\cal E}=40*2*200*10^{-4}*\frac{\pi/3}{0.2}=8.4\:\rm V"